OpenStudy (anonymous):

solve the equation using the quadratic formula. if necessary, round to the nearest hundredth. put in standard form first. x^2+4x=1

sam (.sam.):

\[\Huge Ax^2+Bx+C=0\] Quadratic formula: \[\Huge x=\frac{-B \pm \sqrt{B^2-4AC}}{2A}\] Substitute in.